# Investigating Annotator Bias in Large Language Models for Hate Speech Detection

### v4

@@ 1 Introduction
Within the multifaceted domain of machine learning and Natural Language Processing (NLP), data annotation emerges as a critical juncture, extending beyond simple label assignment to encompass a wide array of supplementary predictive data. This intricate procedure entails a series of nuanced steps: initially, sorting raw data with class or task labels for primary classification, followed by the inclusion of intermediate labels to enrich contextual understanding; subsequently, assigning confidence scores to gauge the reliability of annotations; further, incorporating alignment or preference labels to tailor outputs to specific criteria or user needs; then, annotating entity relationships to illuminate interactions within datasets; delineating semantic roles to uncover the underlying functions of entities within sentences; and finally, tagging temporal sequences to capture the chronological flow of events or actions.
Data annotation poses notable challenges for contemporary machine learning models, owing to the complexity, subjectivity, and diversity of the data involved. It demands domain expertise and entails labor-intensive manual labeling of extensive datasets. Additionally, the process of annotating data is often labor-intensive, requiring significant human effort and expertise. In recent years, the emergence of LLMs, such as OpenAI’s GPT series Radford et al. (2019); Brown et al. (2020) and Google’s BERT Devlin et al. (2018), has revolutionized natural language processing tasks by demonstrating remarkable capabilities in understanding and generating human-like text.
@@ 2 Related Work
The advent of LLMs has revolutionized NLP tasks by enabling the development of more sophisticated and context-aware language understanding systems. Models such as BERT (Devlin et al., 2018), GPT (Radford et al., 2018), and their variants have demonstrated remarkable performance across a wide range of NLP tasks, including text classification, language generation, and question answering. These models leverage pre-training on large corpora followed by fine-tuning on task-specific data, allowing them to capture intricate linguistic patterns and semantic relationships.
@@
The investigation of social biases within Natural Language Processing (NLP) models constitutes a significant area of research. Previous studies have delineated two primary categories of biases and harms: allocational harms and representational harms (Blodgett et al., 2020; Crawford, 2017). Scholars have explored various methodologies to assess and alleviate these biases in both Natural Language Understanding (NLU) (Bolukbasi et al., 2016; Dixon et al., 2018; Zhao et al., 2018; Bordia and Bowman, 2019; Dev et al., 2021; Sun and Peng, 2021) and Natural Language Generation (NLG) tasks (Sheng et al., 2019; Dinan et al., 2019).
Within this body of literature, (Sun and Peng, 2021) proposed utilizing the Odds Ratio (OR) (Szumilas, 2010) as a metric to quantify gender biases, particularly in items exhibiting significant frequency disparities or high salience among genders. (Sheng et al., 2019) assessed biases in NLG model outputs conditioned on specific contextual cues, while (Dhamala et al., 2021) extended this analysis by incorporating real-world prompts extracted from Wikipedia. Several strategies (Sheng et al., 2020; Liu et al., 2021; Cao et al., 2022; Gupta et al., 2022) have been proposed to mitigate biases in NLG models, yet their applicability to closed API-based LLMs, such as ChatGPT, remains uncertain.
@@ 3.1 Data Collection and Annotation
Acknowledging the inherent vagueness in prior methodologies for annotating hate speech, as noted by (Schmidt and Wiegand, 2017), which often led to low agreement scores, our study took measures to enhance the clarity and consistency of the an- notation process. To achieve this, all annotators collaboratively formulated and refined annotation guidelines to ensure a shared understanding of hate speech. An explicit definition, accompanied by a detailed explanation, was provided to elucidate the concept further.

### v5

@@ 1 Introduction
The growing widespread presence of online hate speech presents a critical challenge for maintaining safe and inclusive digital environments. Automated hate speech detection systems, powered by machine learning and large language models (LLMs), have emerged as vital tools to address this issue (Tan et al., 2024). These systems rely heavily on annotated datasets to train and evaluate their performance. However, the process of annotating hate speech is inherently subjective and influenced by the annotators’ sociocultural, and personal biases. As a result, these biases can inadvertently be embedded into the datasets and, subsequently, into the detection models, raising concerns about fairness, accuracy, and generalizability. Despite the advancements in LLMs and their capabilities, annotator bias remains an underexplored yet significant factor affecting their performance and reliability in hate speech detection tasks.
@@ 2 Related Work
The advent of LLMs has revolutionized NLP tasks by enabling the development of more sophisticated and context-aware language understanding systems. Models such as BERT (Devlin et al., 2018), GPT, and their variants have demonstrated remarkable performance across a wide range of NLP tasks, including text classification, language generation, and question answering. These models leverage pre-training on large corpora followed by fine-tuning on task-specific data, allowing them to capture intricate linguistic patterns and semantic relationships.
@@
The investigation of social biases within Natural Language Processing (NLP) models constitutes a significant area of research. Previous studies have delineated two primary categories of biases and harms: allocational harms and representational harms (Blodgett et al., 2020; Crawford, 2017). Scholars have explored various methodologies to assess and alleviate these biases in both Natural Language Understanding (NLU) (Bolukbasi et al., 2016; Dixon et al., 2018; Zhao et al., 2018; Bordia & Bowman, 2019; Dev et al., 2021; Sun & Peng, 2021) and Natural Language Generation (NLG) tasks (Sheng et al., 2019; Dinan et al., 2019).
Within this body of literature, (Sun & Peng, 2021) proposed utilizing the Odds Ratio (OR) (Szumilas, 2010) as a metric to quantify gender biases, particularly in items exhibiting significant frequency disparities or high salience among genders. (Sheng et al., 2019) assessed biases in NLG model outputs conditioned on specific contextual cues, while (Dhamala et al., 2021) extended this analysis by incorporating real-world prompts extracted from Wikipedia. Several strategies (Sheng et al., 2020; Liu et al., 2021; Cao et al., 2022; Gupta et al., 2022) have been proposed to mitigate biases in NLG models, yet their applicability to closed API-based LLMs, such as ChatGPT, remains uncertain.
@@ 3.1 Data Collection and Annotation
Acknowledging the inherent vagueness in prior methodologies for annotating hate speech, as noted by (Schmidt & Wiegand, 2017), which often led to low agreement scores, our study took measures to enhance the clarity and consistency of the an- notation process. To achieve this, all annotators collaboratively formulated and refined annotation guidelines to ensure a shared understanding of hate speech. An explicit definition, accompanied by a detailed explanation, was provided to elucidate the concept further.